OpenStudy (anonymous):

Find a quadratic equation with roots -1 + 4i and -1 - 4i

OpenStudy (anonymous):

Can someone please explain this to me step by step im so lost :(

OpenStudy (ranga):

If p and q are the roots of a quadratic equation, then the quadratic equation will be: (x - p)(x - q) = 0

OpenStudy (ranga):

Therefore, multiply and set it to zero: { x - (-1 + 4i) } * { x - (-1 - 4i) } = 0 Simplify. Remember i^2 = -1

OpenStudy (anonymous):

Im so lost:(

OpenStudy (anonymous):

I need someone to show me this step by step i haven't done this in a really long time so i dont remember anything im so sorry

OpenStudy (ranga):

I typed a whole bunch of stuff and the page refreshed automatically and I lost everything :(

OpenStudy (ranga):

I will type it in Notepad and copy and paste it here. Give me a couple of minutes.

OpenStudy (anonymous):

Okay i appreciate this very much <3 so sorry :(

OpenStudy (ranga):

{ x - (-1 + 4i) } * { x - (-1 - 4i) } Get rid of the inside parenthesis by distributing the negative sign (x + 1 - 4i) * (x + 1 + 4i) Take the first term x on the left and multiply each term on the right with it: x^2 + x + 4xi Take the second term 1 on the left and multiply each term on the right with it: x + 1 + 4i Take the third term -4i on the left and multiply each term on the right with it: -4xi - 4i - 16i^2 = -4xi - 4i + 16 (because i^2 = -1) Add up all three products: x^2 + x + 4xi + x + 1 + 4i + -4xi - 4i + 16 simplify x^2 + 2x + 17 = 0 is the quadratic equation.

OpenStudy (anonymous):

And that is the answer? Btw. Thank you so much you are a life saver :)